Find the slope of an assumed linear demand curve for theater tickets, when persons purchase 1,000 at $5.00 per ticket and 200 at $15.00 per ticket.



1
Expert's answer
2021-07-11T19:40:01-0400

Slope of linear demand curve"=\frac{price\space change}{quantity\space change}=\frac{5-15}{1000-200}=\frac{-10}{800}=\frac{-1}{80}"